OpenStudy (anonymous):

Need some help please. Express F(t) in terms of the alpha function and find L{F(t)}. F(t) = 3 01

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@zepdrix @nincompoop

OpenStudy (anonymous):

Is the "alpha function" defined as \[\alpha(x-c)=\begin{cases}1&\text{for }x\ge c\\0&\text{for }x<c\end{cases}~~?\] If it is, it's more commonly known as the Heaviside theta/step function. And you're supposed to find the Laplace transform of the given function, correct?

OpenStudy (anonymous):

Yes... I was actually able to figure it out.

OpenStudy (anonymous):

Its not bad, I just needed to get the hang of it and I got it \